OpenStudy (anonymous):

The area of a rectangular field is given by the trinomial t^2 – 4t – 45. The length of the rectangle is t + 5. What is the expression for the width of the field?

OpenStudy (anonymous):

t-9

OpenStudy (anonymous):

t^2 – 4t – 45= (t + 5)(t-9) so length is = t-9

OpenStudy (anonymous):

(t+5)(t-9) = t^2-4t-45 l